# Many-Body Localization in the Age of Classical Computing

### v2

@@ XII.2 Outlook
Presently, the status of the MBL field resembles, in some aspects, the situation of glassy systems. In spin glass models, the mean-field picture is understood through the mechanism of replica symmetry breaking [469]. The value of the upper critical dimension, i.e., the spatial dimension above which the mean-field solution applies, however is not clear. Therefore, it is debated whether the mean-field picture explains the results in three dimensions better than the alternative phenomenology of droplets [408, 86, 206] [for a recent review of this debate see [471, 426] and references therein]. In configurational glasses, as reviewed by [63], simulations and experiments show very slow dynamics, and there is no clear evidence. Still, characterization of the microscopic mechanisms responsible for the different physical regimes in configurational glasses remains an important challenge. Similarly, studies of the microscopic mechanisms associated with the interplay of ergodicity and ergodicity breaking in the vicinity of the putative MBL transition may unravel novel physical phenomena, essential for our understanding of non-equilibrium physics of disordered many-body systems, irrespective of the fate of the ETH-MBL crossover in the asymptotic limit.
@@ XIV Appendix: different kinds of transport and compatible equilibria
When discussing localization, and in particular MBL, two clearly distinct but related phenomena are involved: transport (i.e. dynamical relaxation from an initial condition or non-equilibrium steady states) and equilibrium, either in terms of equilibrium distributions or, in more modern terms, eigenstates properties. Can one have different kinds of transport and the same equilibrium state? [^41] The viceversa is probably easily excluded. In principle, the answer is yes. While it is natural to associate diffusion to the Gibbs distribution, a diffusion coefficient to a temperature and a mobility (like in Einstein’s relation) it is possible to construct counterexamples of classsical dynamical processes with anomalous transport but which nonetheless converge to Gibbs equilibrium distribution. So, if the question is: from the mere observation of ETH in the eigenstates, can we conclude that there is a non-zero conductivity/diffusion coefficient in the system, the answer is no. Let us give a couple of examples to show how this works.
We follow closely [327], Chapter 5. Consider a continuous time random walk with independent random time and space jumps $\lambda(\xi)$ and $\psi(\tau)$. The dispersion relation in general connects the typical values of the time passed and the space covered:
@@
where $x=\sum_{i=1}^{n}\xi_{i}$ where $n$ is such that $\sum_{i=1}^{n}\tau_{n}=t$.
For $\gamma=1/2$ we have diffusion, $\gamma<1/2$ is sub-diffusion. A more complete description is contained in the probability distribution $P(x,t)$ or equivalently its Fourier transform $P(k,t)$ or the Fourier-Laplace transform $P(k,s)$.
The same $\gamma$ can be obtained in two ways: a) the waiting time $\psi(\tau)\sim 1/\tau^{1+\alpha}$, for $\alpha<1$ this distribution is long tailed (such that the average $\langle\tau\rangle$ diverges). The distribution $\psi(\xi)$ is Gaussian (or has at least two finite moments). In this case
$P(k,t)\sim E_{\alpha}(-k^{2}t^{\alpha})$ (95)
where $E_{\alpha}(x)=\sum_{n\geq 0}(-x)^{n}/\Gamma(1+\alpha n)$. We have $\langle x^{2}\rangle=t^{\alpha}$ so $\gamma=\alpha/2$. In the Fokker-Plank equation this corresponds to a fractional derivative in time
@@
When a force $F(x)=-\partial V/\partial x$ is introduced, this modifies the right-hand side of the equation
@@
The stationary solution is however still of the Gibbs form: $P\propto e^{-\beta V}$ with $\beta$ related to $D,\mu$ by Einstein relation.
@@
$\frac{\partial}{\partial t}P=-\mu\frac{\partial}{\partial x}(FP)+D_{\gamma}\frac{\partial^{2\gamma}}{\partial x^{2\gamma}}P.$ (98)
where $\gamma>1$ (resp. $<1$) corresponds to subdiffusion (resp. super-diffusion). The equilibrium distribution is not Gibbs anymore, since $e^{-\beta V}$ does not solve the equation with $\partial P/\partial t=0$. It must be mentioned, however, that the case with $\gamma<1$, which corresponds to Levi flights, is more natural than the case $\gamma>1$ since the latter requires fine-tuning to eliminate the $\partial^{2}/\partial x^{2}$ terms which arise naturally and are dominant with respect to $\partial^{2\gamma}/{\partial x^{2\gamma}}$ in a long-wavelength expansion.
Summarizing, subdiffusion is compatible with the Boltzmann-Gibbs distribution, and although this is a statement in classical mechanics, it is most likely the case that subdiffusion is compatible with ETH as well, which explains why it is observed in transport in quantum spin chains at very small disorder [673]. Superdiffusion, instead, leads to non-Gibbsian equilibrium distribution, solution of the fractional derivative equation

### v3

@@ XII.2 Outlook
Presently, the status of the MBL field resembles, in some aspects, the situation of glassy systems. In spin glass models, the mean-field picture is understood through the mechanism of replica symmetry breaking [507]. The value of the upper critical dimension, i.e., the spatial dimension above which the mean-field solution applies, however is not clear. Therefore, it is debated whether the mean-field picture explains the results in three dimensions better than the alternative phenomenology of droplets [444, 89, 219] [for a recent review of this debate see [509, 462] and references therein]. In configurational glasses, as reviewed by [66], simulations and experiments show very slow dynamics, and there is no clear evidence for either of the scenarios. Still, characterization of the microscopic mechanisms responsible for the different physical regimes in configurational glasses remains an important challenge. Similarly, studies of the microscopic mechanisms associated with the interplay of ergodicity and ergodicity breaking in the vicinity of the putative MBL transition may unravel novel physical phenomena, essential for our understanding of non-equilibrium physics of disordered many-body systems, irrespective of the fate of the ETH-MBL crossover in the asymptotic limit.
@@ XIV Appendix: different kinds of transport and compatible equilibria
When discussing localization, and in particular MBL, two clearly distinct but related phenomena are involved: transport (i.e., dynamical relaxation from an initial condition or non-equilibrium steady states) and equilibrium, either in terms of equilibrium distributions or, in more modern terms, eigenstate properties. Can one have different kinds of transport and the same equilibrium state? [^51] The vice versa is probably easily excluded. In principle, the answer is yes. While it is natural to associate diffusion to the Gibbs distribution, a diffusion coefficient to a temperature and mobility (like in Einstein’s relation), it is possible to construct counterexamples of classical dynamical processes with anomalous transport but which nonetheless converge to Gibbs equilibrium distribution. So, if the question is: from the mere observation of ETH in the eigenstates, can we conclude that there is a non-zero conductivity/diffusion coefficient in the system, the answer is no. Let us give a couple of examples to show how this works.
We follow closely [354], Chapter 5. Consider a continuous time random walk with independent random time and space jumps $\lambda(\xi)$ and $\psi(\tau)$. The dispersion relation in general connects the typical values of the time passed and the space covered,
@@
where $x=\sum_{i=1}^{n}\xi_{i}$ and $n$ is such that $\sum_{i=1}^{n}\tau_{n}=t$.
For $\gamma=1/2$, we have diffusion, and $\gamma<1/2$ is sub-diffusion. A more complete description is contained in the probability distribution $P(x,t)$ or equivalently its Fourier transform $P(k,t)$, or the Fourier-Laplace transform $P(k,s)$.
The same $\gamma$ can be obtained in two ways.
a) The waiting time $\psi(\tau)\sim 1/\tau^{1+\alpha}$, for $\alpha<1$, this distribution is long tailed (such that the average $\langle\tau\rangle$ diverges). The distribution $\psi(\xi)$ is Gaussian (or has at least two finite moments). In this case,
$P(k,t)\sim E_{\alpha}(-k^{2}t^{\alpha}),$ (95)
where $E_{\alpha}(x)=\sum_{n\geq 0}(-x)^{n}/\Gamma(1+\alpha n)$. We have $\langle x^{2}\rangle=t^{\alpha}$, so $\gamma=\alpha/2$. In the Fokker-Plank equation, this corresponds to a fractional derivative in time,
@@
When a force $F(x)=-\partial V/\partial x$ is introduced, this modifies the right-hand side of the equation,
@@
The stationary solution is however still of the Gibbs form, $P\propto e^{-\beta V}$, with $\beta$ related to $D,\mu$ by the Einstein relation.
@@
$\frac{\partial}{\partial t}P=-\mu\frac{\partial}{\partial x}(FP)+D_{\gamma}\frac{\partial^{2\gamma}}{\partial x^{2\gamma}}P,$ (98)
where $\gamma>1$ ($\gamma<1$) corresponds to subdiffusion (superdiffusion). The equilibrium distribution is not Gibbs anymore, since $e^{-\beta V}$ does not solve the equation with $\partial P/\partial t=0$. It must be mentioned, however, that the case with $\gamma<1$, which corresponds to Levi flights, is more natural than the case $\gamma>1$ since the latter requires fine-tuning to eliminate the $\partial^{2}/\partial x^{2}$ terms, which arise naturally and are dominant with respect to $\partial^{2\gamma}/{\partial x^{2\gamma}}$ in a long-wavelength expansion.
Summarizing, subdiffusion is compatible with the Boltzmann-Gibbs distribution, and although this is a statement in classical mechanics, it is most likely the case that subdiffusion is compatible with some version of the ETH as well, which explains why it is observed in transport in quantum spin chains at very small disorder [727]. Superdiffusion, instead, leads to non-Gibbsian equilibrium distribution, solution of the fractional derivative equation